OpenStudy (anonymous):

Evaluate the integral by changing to spherical coordinates: -3 sqrt(9-x^2) sqrt(9-x^2-y^2) ∫ ∫ ∫ zsqrt(x^2+y^2+z^2)dzdydx 3 -sqrt(9-x^2) 0 Please explain step by step

OpenStudy (anonymous):

Please typeset your question so it would be clear which intgral and waht function to deal with. Thanks

OpenStudy (anonymous):

\[\int\limits\limits_{-3}^{3}\int\limits_{-\sqrt{9-x^2}}^{\sqrt{9-x^2}}\int\limits_{0}^{\sqrt{9-x^2-y^2}}z \sqrt{x^2+y^2+z^2}dzdydx\]

OpenStudy (anonymous):

i got it on my own

OpenStudy (anonymous):

@Janex, so you do not need any help for that problem.

OpenStudy (anonymous):

@eliassaab no, but thanks!